You lose a coin from your purse. You know that your purse held 80 pennies and 10 dimes. What is the probability that the lost coin is a penny? What is the complement -that the lost coin is NOT a penny?
Answer by ewatrrr(24785)

90 COINS: 10d, 80P
Lost one
P( PENNY) = 80/90 = 8/9
P(not a Penny) = P(Dime) = 10/90 = 1/9